# Quantifying interdisciplinary synergy in higher STEM education

### v1

@@ 1 Introduction
The primary objective of this research is to establish a principled data-driven framework for designing and evaluating interdisciplinary-integrated curricula. To achieve this, we analyze large-scale real-world data extending beyond traditional curriculum design approaches. We develop a data process method and quantifying of interdisciplinary. Our methodology begins with standardizing extensive scientific and technological natural language data. We build a standardized dataset using cutting-edge natural language processing (NLP) techniques, such as large-scale language models (LLM) and higher-order methods. We link this dataset to departments across institutions through a bipartite network to construct knowledge maps. By projecting this network, we analyze standard subjects’ relationships and interconnections among various disciplines.
In particular, we focus on the science, technology, engineering, and math (STEM) disciplines. The decision to concentrate on the STEM disciplines in our analysis is largely influenced by the dynamic evolution and heightened demand these disciplines are experiencing, driven by recent technological advancements. The STEM disciplines are at the forefront of innovation, playing a crucial role in addressing some of the world’s most pressing challenges. As a result, there is an increasing demand for professionals skilled in these areas, necessitating a stronger emphasis on interdisciplinarity to equip future experts with the ability to integrate knowledge and skills from various domains [7]. New policies are advocating for an interdisciplinary approach within STEM education, recognizing that the complexity of modern challenges requires solutions that are not confined to a single discipline. By focusing on STEM, this research aims to contribute to the development of educational frameworks that not only respond to these policy shifts but also prepare students to thrive in a rapidly evolving technological landscape.
@@ 2 Data processing of course standardization
As introduced in Sec. 1, we focus on the departments running four-year undergraduate programs within STEM fields classified under Engineering and Natural Sciences in the Standard Classification (the highest-level classification). Specifically, within Natural Sciences, we extracted data from departments categorized as “Basic Natural Sciences,” which include “Chemistry, Life Science, and Environment” and “Mathematics, Physics, Astronomy, and Earth Science” in the Standard Classification (the intermediate-level classification). The outcome of preprocessing and our comprehensive analysis encompasses curricula from 2 84128412\,8412 841 STEM departments [^1] The original data included 2,851 unique STEM undergraduate departments. However, due to military security restrictions and data structure errors, only 2,841 departments were ultimately analyzed. across $161$ higher education institutions, involving 126 437126437126\,437126 437 “raw” courses.
@@ 2.2 Constructing hyperedges among raw courses with LLM
Now we have the set of raw courses for each token $t$ as in Fig. 1(b), and our next step is to collect the effectively same raw courses as intermediate-level units. Mathematically, it corresponds to identifying higher-order relationships (or hyperedges) among the raw courses than pairwise ones that reflect semantic disciplinary characteristics [13, 14, 15, 16]. Similar to the tokenization process described in Sec. 2.1, we again use a prompt that effectively clusters identical courses based on shared topics or subfields (prompt 2 in Appendix A), depicted in Fig. 1(b). Through this higher-order clustering process, we generate 91 9649196491\,96491 964 hyperedges from all tokens, which encapsulate the intermediate-level interrelationships present in the raw course data.
@@ 4.1 Single-discipline curriculum model
To decide if we accept a candidate course in the model, we calculate the Jensen-Shannon divergence (JSD) between the model weight vector $\mathbf{w}_{\mathrm{model}}$ (the model curriculum vector when the candidate course is included) and the set of real credit vectors $\mathbf{w}_{d}$ of department $d$ belonging to discipline $\mathcal{D}$. More specifically, we quantify the level of congruence of a candidate course with the JSD,
@@
where $\tilde{w}_{\mathcal{D};i}$ is the credit of course $i$ in the optimized curriculum of discipline $\mathcal{D}$.
@@ 4.3 Quantifying interdisciplinarity with the curriculum synergy score
In Sec. 4.2, we have proposed an information-theoretical model for interdisciplinary curricula. Its object of optimization is to design the most congruent and realistic set of courses covering both disciplines. For a systematic approach, we introduce the curriculum synergy score (CSS) based on pedagogy and information theory. As the first step, we quantify the synergy of our top-down interdisciplinary curricula $\Omega\left(\mathcal{D}_{A}\otimes\mathcal{D}_{B}\right)$, compared with a simple bottom-up double major model by hypothetical students denoted by $\Omega\left(\mathcal{D}_{A}\cup\mathcal{D}_{B}\right)$. Although the concept of synergy is nontrivial, we suggest a measure incorporating two key aspects: (1) there should be a decent level of similarity between the two disciplines for an interdisciplinary curriculum, and (2) there should be enough amount of new information (“surprisal” in information theory [8]) emerged from the interdisciplinary curriculum, with respect to simple course-taking activities for the double-major. We assume the following course-weighted vector for the double-major,
@@
Let us explore the mathematical interpretation of integrated disciplines through the lens of pedagogical theory. In contrast to multidisciplinarity, where disciplines coexist independently without significant integration, interdisciplinarity focuses on the meaningful synthesis of knowledge, tools, and perspectives from multiple fields [24, 25]. Achieving interdisciplinary synergy requires a substantial overlap between disciplines, as opposed to merely combining them in a multidisciplinary manner. Our JSD-based CSS is specifically designed to capture this distinction by quantifying how effectively an integrated curriculum encompasses both shared similarities and the emergence of new information, as compared to a basic multidisciplinary curriculum. To quantify the synergy reflecting those aspects, we employ an information-theoretic method based on prior research in educational fields [4, 5, 26, 27, 28, 29]. For the new information emerging from the interdisciplinary curriculum, we again apply the JSD for the course-weighted vectors to quantify the informational distance between $\Omega\left(\mathcal{D}_{A}\cup\mathcal{D}_{B}\right)$ and $\Omega\left(\mathcal{D}_{A}\otimes\mathcal{D}_{B}\right)$, given by
@@
The CSS, therefore, quantifies the synergy between disciplines that cannot solely be captured by simple similarity measures, which highlights the added value of proper interdisciplinary integration. We observe the distribution of each factor separately, and one can check that the vast majority of $\sigma_{\mathcal{D}_{A},\mathcal{D}_{B}}$ values are close to 00 [Fig. 5(a)], which is understandable by the fact that most discipline pairs would be quite different to each other. On the other hand, the distribution of $\mathcal{J}\left(\tilde{\mathbf{w}}_{\mathcal{D}_{A}\cup\mathcal{D}_{B}},\tilde{\mathbf{w}}_{\mathcal{D}_{A}\otimes\mathcal{D}_{B}}\right)$ shows the most prominent peak around $0.3$ presumably corresponding to the overall scale difference between $\tilde{\mathbf{w}}_{\mathcal{D}_{A}\cup\mathcal{D}_{B}}$ and $\tilde{\mathbf{w}}_{\mathcal{D}_{A}\otimes\mathcal{D}_{B}}$, and the second tallest peak around $0.6$ [Fig. 5(b)]. There is no notable correlation between $\sigma_{\mathcal{D}_{A},\mathcal{D}_{B}}$ and $\mathcal{J}\left(\tilde{\mathbf{w}}_{\mathcal{D}_{A}\cup\mathcal{D}_{B}},\tilde{\mathbf{w}}_{\mathcal{D}_{A}\otimes\mathcal{D}_{B}}\right)$ [Fig. 5(c)], which also justifies the formalism of $\Theta\left(\mathcal{D}_{A},\mathcal{D}_{B}\right)$ as the product of them. For notable examples, we provide three examples in supplementary tables: two examples of beneficial synergy (‘physics’ + ‘semiconductor engineering’ in Table S1. and ‘chemistry’ + ‘high polymer engineering’ in Table S2.) and an example of no synergy (‘physics’ + ‘chemistry’ in Table S3.). It is interesting to note that science disciplines and their related engineering disciplines are much better combinations than two science disciplines, and this tendency will be highlighted again in the analysis presented in the next subsection as well.
@@ 4.4 The CSS network analysis
One might perceive this result as unintuitive considering the role of basic science and engineering, but in terms of academic curricula, the application side (engineering) is more synergetically attached to other disciplines than basic science that endows most fundamental concepts and frameworks as basic resources of understanding the entire STEM fields. In other words, the interdisciplinary synergy is not readily obtained due to the unique role of natural science disciplines, which in turn suggests the necessity of maintaining their established curricula as anchoring sources of STEM education, as suggested by the CSS value of zero between physics and chemistry in Table S3.. In the case of research collaboration [33], of course, the pattern of interdisciplinarity can be different, so more comprehensive understanding across education and research can be a natural candidate for future works.
@@ 5 Discussion and conclusion
The result from the curriculum-wise organizational structure of departments suggests a deeper division between PSE and MLSE than that between science and engineering, with notable reorganization of less traditional departments depending on the organizational scale. The departments belonging to MLSE are more coherently organized, reflected by their intact community structure at finer resolution, while engineering departments show more flexible relationships within STEM disciplines. Such organizational flexibility of engineering departments is also highlighted in the synergetic relationship quantified by the CSS. The engineering disciplines constitute multiple core structures in the CSS network, while basic science disciplines are located in the peripheral part with their unique sources indispensable for STEM education.
Considering the result of our data analysis, we call for a more careful approach to designing interdisciplinary education policy. For instance, the fact that basic science disciplines provide fundamental knowledge to understand STEM does not automatically imply that it is beneficial to merge them with other disciplines; rather, we may need to acknowledge their own merit as fundamental sources of knowledge and skills. As we have demonstrated that this type of quantitative analysis can bring such insights, a natural extension of this work can be the one including non-STEM disciplines such as literature, art, history, etc., possibly across different countries. In particular, in this era of technological breakthroughs and innovations in education, the systematic analysis of temporal change in curriculum structures is also highly desirable. On a theoretical side, the interdisciplinary curriculum models involving three or more related disciplines are a candidate for future work. We hope that our work, equipped with nationwide exhaustive data and LLM, provides a nice example for this direction.
@@ Acknowledgements
Author contributions
Funding
This research was supported by the National Research Foundation of Korea (NRF) under the grants NRF-2021R1C1C1004132 (GG and SHL) and NRF-2022R1A2C1091324 (JY).
Data availability
@@ Appendix B Example of tokenization and course standardization
$C_{4}\equiv\text{Statistical Thermodynamics:}$: $\quad T(C_{4})=\{\text{Statistics},\text{Thermo},\text{Dynamics}\}.$
@@
$\mathcal{G}(\text{Statistics})=\{C_{1},C_{2},C_{4}\},\quad\mathcal{G}(\text{Thermo})=\{C_{3},C_{4}\}.$
@@
  • For the token Statistics,
    $\mathcal{G}(\text{Statistics})=\{C_{1},C_{2},C_{4}\}\,,$
    and the LLM determines that $C_{1}$ and $C_{4}$ are semantically similar, forming a hyperedge,
    $H_{1}=\{C_{1},C_{4}\}\,.$
@@
    $\mathcal{G}(\text{Thermo})=\{C_{3},C_{4}\}\,.$
@@
Using the hyperedges $H_{1}$ and $H_{2}$, we construct a hypergraph. The hypergraph contains:
$H_{1}=\{C_{1},C_{4}\},\quad H_{2}=\{C_{3},C_{4}\}.$
@@
$\mathcal{C}_{1}=\{C_{1},C_{3},C_{4}\}\,.$
Therefore, the raw courses $C_{1}$: Statistical Physics, $C_{3}$: Thermodynamics, and $C_{4}$: Statistical Thermodynamics are grouped into a single standardized course, in regard to statistical physics.

### v2

@@ 1 Introduction
The primary objective of this research is to establish a principled data-driven framework for designing and evaluating interdisciplinary-integrated curricula. To achieve this, we analyze large-scale real-world data extending beyond traditional curriculum design approaches. We develop a principled data processing pipeline and a quantitative framework to measure interdisciplinarity across large-scale academic curricula. Our methodology begins with standardizing extensive scientific and technological natural language data. We build a standardized dataset using cutting-edge natural language processing (NLP) techniques, such as large-scale language models (LLMs) [7], and higher-order methods [8]. We link this dataset to departments across institutions through a bipartite network to construct knowledge maps. By projecting this network, we analyze standard subjects’ relationships and interconnections among various disciplines.
In particular, we focus on the science, technology, engineering, and math (STEM) disciplines. The decision to concentrate on the STEM disciplines in our analysis is largely influenced by the dynamic evolution and heightened demand these disciplines are experiencing, driven by recent technological advancements. The STEM disciplines are at the forefront of innovation, playing a crucial role in addressing some of the world’s most pressing challenges. As a result, there is an increasing demand for professionals skilled in these areas, necessitating a stronger emphasis on interdisciplinarity to equip future experts with the ability to integrate knowledge and skills from various domains [9]. Recent policy directions, including those led by the U.S. National Science Foundation, emphasize the necessity of interdisciplinary training in STEM to prepare the next generation of leaders for complex, real-world challenges [10]. New policies are advocating for an interdisciplinary approach within STEM education, recognizing that the complexity of modern challenges requires solutions that are not confined to a single discipline [11]. By focusing on STEM, this research aims to contribute to the development of educational frameworks that not only respond to these policy shifts but also prepare students to thrive in a rapidly evolving technological landscape.
@@ 2 Data processing of course standardization
As introduced in Sec. 1, we focus on the departments running four-year undergraduate programs within STEM fields classified under Engineering and Natural Sciences in the Standard Classification (the highest-level classification). Specifically, within Natural Sciences, we extracted data from departments categorized as “Basic Natural Sciences,” which include “Chemistry, Life Science, and Environment” and “Mathematics, Physics, Astronomy, and Earth Science” in the Standard Classification (the intermediate-level classification). As a result of our preprocessing and comprehensive data integration, we constructed a standardized dataset comprising curricula from 2 84128412\,8412 841 STEM departments [^1] The original data included 2 85128512\,8512 851 unique STEM undergraduate departments. However, due to military security restrictions and data structure errors, only 2 84128412\,8412 841 departments were ultimately analyzed. across $161$ higher education institutions, including a total of 126 437126437126\,437126 437 raw course entries. We use departments as the primary unit of analysis because our data, sourced from the Ministry of Education in Korea, is organized at this level, providing a consistent framework with around $80$ credits (for major courses) over four years per department. In Korea, departments serve as the de facto units of disciplines due to the structured nature of higher education, aligning with current educational practices and government initiatives that encourage interdisciplinary integration. While finer distinctions at the program level might be desirable, current data limitations necessitate this approach. Future research could explore more granular analyses as detailed data becomes available, such as through “microdegree” programs aimed at fostering small units of interdisciplinary education.
We also note that different departments may offer courses with varying credit allocations, but the overall credit requirement standardizes the instructional volume, facilitating consistent cross-departmental comparisons. All curricula are assessed under a rather uniform constraint in the total number of credits, ensuring that variations in course numbers do not bias our analysis. This approach is again consistent with Korea’s centralized higher education policy, characterized by a national standard for four-year undergraduate curricula. Although such discrepancies might have greater implications in countries with more flexible curricula, their impact would be minimal in the Korean educational context.
@@ 2.2 Constructing hyperedges among raw courses with LLM
Now we have the set of raw courses for each token $t$ as in Fig. 1(b), and our next step is to collect the effectively same raw courses as intermediate-level units. Mathematically, this step involves identifying higher-order relationships (i.e., hyperedges) [8] among raw courses—beyond simple pairwise associations—so as to capture more nuanced semantic groupings that reflect disciplinary structures [16, 17, 18]. Similar to the tokenization process described in Sec. 2.1, we again use a prompt that effectively clusters identical courses based on shared topics or subfields (prompt 2 in Appendix A), depicted in Fig. 1(b).
While conventional NLP techniques have long been applied to large-scale textual data analysis, they often rely on rigid pipelines tailored to narrow tasks. In contrast, we adopt large language models (LLMs) due to their superior contextual understanding and adaptability. Recent studies have demonstrated that LLMs consistently outperform traditional NLP tools in extracting nuanced semantic relationships from domain-specific scientific corpora [19]. Their ability to integrate context, handle ambiguous terminology, and operate with minimal labeled data makes them especially effective for tasks involving the clustering of STEM course descriptions, which are rich in discipline-specific language and complex structural patterns.
Through this LLM-based higher-order clustering process, we generate 91 9649196491\,96491 964 hyperedges from all tokens, which encapsulate the intermediate-level interrelationships present in the raw course data.
@@ 4.1 Single-discipline curriculum model
To evaluate the degree of new information introduced by the interdisciplinary curriculum, we compute the Jensen-Shannon divergence (JSD) between the course-weighted vectors of the simple union and the integrated model. We denote this divergence as $\mathcal{J}$, which provides a symmetric and bounded measure of dissimilarity between the two distributions:
@@
where $\tilde{w}_{\mathcal{D};i}$ is the credit of course $i$ in the optimized curriculum of discipline $\mathcal{D}$. The vector for each discipline is standardized by the optimization process described in Sec. 2, resulting in the uniform credit sum $\approx 80$, to ensure comparability and consistency in the curriculum design process.
@@ 4.3 Quantifying interdisciplinarity with the curriculum synergy score
In Sec. 4.2, we have proposed an information-theoretical model to evaluate the degree of integration achieved by interdisciplinary curricula. The central aim is to design a curriculum that reflects both semantic cohesion across departments and pedagogical completeness within a combined field of study. To formalize this, we define an optimized curriculum set $\Omega(\mathcal{D})$ for any given set of departments $\mathcal{D}$. The selection process is based on a semantic filtering criterion: each course must contribute to reducing the average divergence in Eq. (5) across departmental curricula. The credit constraint mentioned in Sec. 4.1 ensures that the resulting curriculum maintains the completeness required for a full major according to national education standards. An analogous formulation is applied to interdisciplinary combinations $\mathcal{D}_{A}\otimes\mathcal{D}_{B}$. Based on this formal construction, we introduce the curriculum synergy score (CSS), which evaluates whether the algorithmically constructed interdisciplinary curriculum $\Omega\left(\mathcal{D}_{A}\otimes\mathcal{D}_{B}\right)$ achieves a deeper level of integration than a simple combination of the two disciplines, denoted by $\Omega\left(\mathcal{D}_{A}\cup\mathcal{D}_{B}\right)$.
Although the concept of synergy is nontrivial, our operationalization focuses on two key aspects: (1) the interdisciplinary curriculum should maintain a sufficient degree of alignment with its constituent disciplines; and (2) it should generate new semantic content that is not trivially inherited from the union of its parts. This second aspect is grounded in the notion of “surprisal” from information theory [12], reflecting the emergence of integrative knowledge structures not found in either discipline alone. We assume the following course-weighted vector for the double-major,
@@
Let us explore the mathematical interpretation of integrated disciplines through the lens of pedagogical theory. In contrast to multidisciplinarity, where disciplines coexist independently without significant integration, interdisciplinarity focuses on the meaningful synthesis of knowledge, tools, and perspectives from multiple fields [27, 28]. Achieving interdisciplinary synergy requires a substantial overlap between disciplines, as opposed to merely combining them in a multidisciplinary manner. Our JSD-based CSS is specifically designed to capture this distinction by quantifying how effectively an integrated curriculum encompasses both shared similarities and the emergence of new information, as compared to a basic multidisciplinary curriculum. This approach helps distinguish truly integrative designs from mere aggregations by evaluating whether the synthesized curriculum provides added pedagogical value rather than duplicated content. For students, a high CSS implies a more coherent learning experience, minimizing redundancy and promoting cross-disciplinary connections that are more than the sum of their parts. To quantify the synergy reflecting those aspects, we employ an information-theoretic method based on prior research in educational fields [4, 5, 29, 30, 31, 32]. For the new information emerging from the interdisciplinary curriculum, we again apply the JSD for the course-weighted vectors to quantify the informational distance between $\Omega\left(\mathcal{D}_{A}\cup\mathcal{D}_{B}\right)$ and $\Omega\left(\mathcal{D}_{A}\otimes\mathcal{D}_{B}\right)$, given by
@@
The CSS, therefore, quantifies the synergy between disciplines that cannot solely be captured by simple similarity measures, which highlights the added value of proper interdisciplinary integration. We observe the distribution of each factor separately, and one can check that the vast majority of $\sigma_{\mathcal{D}_{A},\mathcal{D}_{B}}$ values are close to 00 [Fig. 5(a)], which is understandable by the fact that most discipline pairs would be quite different to each other. On the other hand, the distribution of $\mathcal{J}\left(\tilde{\mathbf{w}}_{\mathcal{D}_{A}\cup\mathcal{D}_{B}},\tilde{\mathbf{w}}_{\mathcal{D}_{A}\otimes\mathcal{D}_{B}}\right)$ shows the most prominent peak around $0.3$ presumably corresponding to the overall scale difference between $\tilde{\mathbf{w}}_{\mathcal{D}_{A}\cup\mathcal{D}_{B}}$ and $\tilde{\mathbf{w}}_{\mathcal{D}_{A}\otimes\mathcal{D}_{B}}$, and the second tallest peak around $0.6$ [Fig. 5(b)]. There is no notable correlation between $\sigma_{\mathcal{D}_{A},\mathcal{D}_{B}}$ and $\mathcal{J}\left(\tilde{\mathbf{w}}_{\mathcal{D}_{A}\cup\mathcal{D}_{B}},\tilde{\mathbf{w}}_{\mathcal{D}_{A}\otimes\mathcal{D}_{B}}\right)$ [Fig. 5(c)], which also justifies the formalism of $\Theta\left(\mathcal{D}_{A},\mathcal{D}_{B}\right)$ as the product of them. In Appendix D, we provide the plots in the double-logarithmic scale in Fig. 9 to help the reader better understand the range and tail behavior of the similarity values.
For notable examples, we provide three examples in supplementary tables: two examples of beneficial synergy (‘physics’ + ‘semiconductor engineering’ in Table S1 and ‘chemistry’ + ‘high polymer engineering’ in Table S2) and an example of no synergy (‘physics’ + ‘chemistry’ in Table S3). It is interesting to note that science disciplines and their related engineering disciplines are much better combinations than two science disciplines, and this tendency will be highlighted again in the analysis presented in the next subsection as well. Other examples include combinations ‘biotechnology + energy engineering,’ which are intuitively synergistic, alongside less conventional yet plausible pairs such as ‘nautical science + civil engineering’ and ‘optical Engineering + semiconductor engineering.’
@@ 4.4 The CSS network analysis
One might perceive this result as unintuitive considering the role of basic science and engineering, but in terms of academic curricula, the application side (engineering) is more synergetically attached to other disciplines than basic science, which endows most fundamental concepts and frameworks as basic resources of understanding the entire STEM fields. In other words, the interdisciplinary synergy is not readily obtained due to the unique role of natural science disciplines, which in turn suggests the necessity of maintaining their established curricula as anchoring sources of STEM education, as suggested by the CSS value of zero between physics and chemistry in Table S3. In the case of research collaboration [36], of course, the pattern of interdisciplinarity can be different, so a more comprehensive understanding across education and research can be a natural candidate for future works.
@@ 5 Discussion and conclusion
The result from the curriculum-wise organizational structure of departments suggests a deeper division between PSE and MLSE than that between science and engineering, with notable reorganization of less traditional departments depending on the organizational scale. The departments belonging to MLSE are more coherently organized, reflected by their intact community structure at finer resolution, while engineering departments show more flexible relationships within STEM disciplines. Such organizational flexibility of engineering departments is also highlighted in the synergetic relationship quantified by the CSS. The engineering disciplines constitute multiple core structures in the CSS network, while basic science disciplines are located in the peripheral part, with their unique sources indispensable for STEM education.
Considering the result of our data analysis, we call for a more careful approach to designing interdisciplinary education policy. For instance, the fact that basic science disciplines provide fundamental knowledge to understand STEM does not automatically imply that it is beneficial to merge them with other disciplines; rather, we may need to acknowledge their own merit as fundamental sources of knowledge and skills. However, it is important to acknowledge that our algorithmic approach alone cannot overcome complex educational challenges. When our data-driven insights are integrated with evolving educational needs, we can design more effective curricula that better prepare students for future demands. As we have demonstrated that this type of quantitative analysis can bring such insights, A natural extension of this work can be one that includes non-STEM disciplines such as literature, art, history, etc., possibly across different countries. In particular, in this era of technological breakthroughs and innovations in education, the systematic analysis of temporal change in curriculum structures is also highly desirable. On a theoretical side, the interdisciplinary curriculum models involving three or more related disciplines are a candidate for future work. We hope that our work, equipped with nationwide exhaustive data and LLM, provides a nice example for this direction.
@@
Declarations
Ethics approval and consent to participate
The data sets utilized in this study have been sourced from Higher Education in Korea (대학알리미 in Korean) [13, 14, 15], a public data service of the Ministry of Education in Korea. This study does not incorporate any personal information.
Consent for publication
All of the authors consent to publication.
Availability of data and material
Competing interests
The authors declare that they have no competing interests.
Funding
This research was supported by the National Research Foundation of Korea (NRF) under the grants RS-2021-NR061247 (GG and SHL) and RS-2022-NR069864 (JY).
Author contributions
@@ Appendix B Example of tokenization and course standardization
$C_{4}\equiv\text{Statistical Thermodynamics:}$: $\quad T(C_{4})=\{\text{Statistics},\text{Thermo},\text{Dynamics}\},$
$C_{5}\equiv\text{Statistical Mechanics:}$: $\quad T(C_{5})=\{\text{Statistics},\text{Mechanics}\}.$
@@
$\mathcal{G}(\text{Statistics})=\{C_{1},C_{2},C_{4},C_{5}\},\quad\mathcal{G}(\text{Thermo})=\{C_{3},C_{4}\}.$
@@
  • For the token Statistics, the LLM determines that $C_{1}$, $C_{4}$, and $C_{5}$ are semantically aligned in their scientific context, forming a hyperedge,
    $H_{1}=\{C_{1},C_{4},C_{5}\}\,.$
@@
    $\mathcal{G}(\text{Thermo})=\{C_{3},C_{4}\}\,,$
@@
Using the hyperedges $H_{1}$ and $H_{2}$, we construct a hypergraph:
$H_{1}=\{C_{1},C_{4},C_{5}\},\quad H_{2}=\{C_{3},C_{4}\}.$
@@
$\mathcal{C}_{1}=\{C_{1},C_{3},C_{4},C_{5}\}\,.$
Therefore, the raw courses $C_{1}$: Statistical Physics, $C_{3}$: Thermodynamics, $C_{4}$: Statistical Thermodynamics, and $C_{5}$: Statistical Mechanics are grouped into a single standardized course in the context of statistical physics and thermodynamics.